OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2). A. y = 3x – 5 B. y = 3x – 13 C. y = 3x – 1 D. y = 3x + 2

OpenStudy (anonymous):

since line is parallel to y = 3x +2 its slope is m=3 slope of required line must be 3 because parallel lines have same slope!! so using point slop form \[\Large y-y _{0} =m(x-x _{0})\] where (x0,y0) are given (5,2) \[\Large y-2=3(x-5)\] \[\Large y-2=3x-15\] \[\Large y=3x-15+2\] \[\Large y=3x-13\] i hope it is clear !

OpenStudy (anonymous):

and yes slope of the above is m=3 and y intercept is -13 !

OpenStudy (anonymous):

@dasDaName did you get it ?

OpenStudy (anonymous):

B. yeah i got it, thanks

OpenStudy (anonymous):

you 're welcome :)